OpenStudy (anonymous):

calculus help...plz explain and answer question...i totally dont know what they are...studying on my own

OpenStudy (anonymous):

Do you understand vectors and parametric equations?

OpenStudy (anonymous):

i know a little abput para,etric equations

OpenStudy (hunus):

The position of the hoop is given by \[\vec{r} = (t-\sin(t))\vec{i}+(1-\cos(t))\vec{j}\ So |v| will be the magnitude of the derivative of r and |a| will be the magnitude of the derivative of v

OpenStudy (anonymous):

ok. so how to solve for first and second answer

OpenStudy (hunus):

Do you know how to take derivatives?

OpenStudy (anonymous):

yes sir...i know how to do everything except last two chapters of calculus

OpenStudy (hunus):

well first you're going to be taking \[\frac{d \vec{r}}{dt}\]

OpenStudy (anonymous):

of what

OpenStudy (hunus):

Of the vector r that you're given. To take a derivative of a vector take the derivative of its components and still leave them attached to the components. So for example the derivative of the vector s with components (3t^2 +2)i + 2j would be ds/dt = (6t)*i + 0j

OpenStudy (hunus):

To take a derivative of a vector take the derivative of its components and still leave them attached to their unit vectors**.

OpenStudy (anonymous):

so is the derivative (-costi)+(sintj)

OpenStudy (hunus):

the i component resolves to (1-cos(t))*i

OpenStudy (anonymous):

i is just x and j is just y...right?

OpenStudy (hunus):

Yup

OpenStudy (anonymous):

so the derivative would be - wait what is t then

OpenStudy (hunus):

t is telling you how x and y changes -- t is time

OpenStudy (anonymous):

ohk...would you give me the derivative of first part (t+sint)i...i will get the rest of the part..its just confusing me

OpenStudy (hunus):

Well when you were taking derivatives as a function of x, you were seeing how things changed when you moved x. Now you're just seeing how they move when you change the time. dr/dt = v = d((t+sint)i+(1-cos(t))j)/dt = (1 - cos(t))i + (sin(t))j

OpenStudy (anonymous):

ohk got it...thx...now i know to set them as zero and then solve but again i and j is confusing me

OpenStudy (hunus):

Well some text books use the letters x and y instead of i and j

OpenStudy (hunus):

Just be sure to write an arrow over them if you do use them

OpenStudy (anonymous):

lemme try (1-cost)i + (sint)j = 0 (1-cost)i = - (sint)j (1-cost)/(sint) = - j / i

OpenStudy (hunus):

You're treating i and j like regular variables. They signify a direction (x and y). 1/i has no meaning if i is a vector. (1-cos(t))i Just says that its velocity in the x direction is 1-cos(t) (slowing down and speeding up) for instance the vector \[\vec{m} = 4\vec{i} + 6\vec{j}\] visually looks like this  Your answer for v will simply be (1 - cos(t))i + (sin(t))j